The level of calcium in the blood of healthy young adults varies with a population mean of about 9.5 milligrams per deciliter. A clinic in rural Guatemala measures the blood calcium level of 180 healthy pregnant women at their first visit for prenatal care. The sample mean is 9.54 and sample standard deviation is 0.4. Is this an indication that the mean calcium level in the population from which these women come is greater than 9.5?

1.) What is the p-value?

please explain in detail, I dont understand how to get p-value for this problem
Answer #1

Degree of freedom = 180-1 = 179

t = 1.342

P value = P(t>1.342)

Either you can use Excel or t-tables to calculate this

Use Excel function

1-T.DIST(A40,179,TRUE)

= 0.0907

Since P value is greater than 0.05 ( assumed Signficance level) , we failed to conclude that mean calcium level in the population from which these women come is greater than 9.5